# Examining the influence of anisotropy on the fundamental mode of nonradial oscillation in quark stars on a complete general relativistic scheme

### v1

@@ 4.3 On the detectability of the fundamental mode signal
1.0-1.0 $577.477$ $1.2926$ $3.2316$
0.5-0.5 $530.746$ $1.2912$ $3.2280$
$400$ $0.0$ $485.668$ $1.2897$ $3.2243$
+0.5+0.5 $443.167$ $1.2874$ $3.2186$
+1.0+1.0 $402.307$ $1.2847$ $3.2119$
1.0-1.0 $187.033$ $1.5409$ $3.8523$
0.5-0.5 $174.550$ $1.5349$ $3.8373$
$600$ $0.0$ $164.209$ $1.5267$ $3.8169$
+0.5+0.5 $157.036$ $1.5150$ $3.7875$
+1.0+1.0 $153.927$ $1.4985$ $3.7462$
Table 2: The damping time $\tau$ for strange quark stars, the energy required to excite the fundamental mode, $E^{LV}_{gw}$ and $E^{ET}_{gw}$, for the signal to be detectable by Advanced LIGO-Virgo and the Einstein Telescope, respectively. For this result, we considered the source at $d\sim 10$ $\rm[\rm kpc]$, a signal-to-noise ratio S/N $\sim 5$, two different central densities, and different values of $\alpha$.
In Table 2, we show the minimum energy required for the fundamental mode to be detected. For this objective, we consider two detectors; the first one is Advanced LIGO-Virgo with a sensitivity of $S_{n}^{1/2}\sim 2\times 10^{-23}\,\mathrm{[\rm Hz]}^{-1/2}$ at $\sim[\rm kHz]$ [1], and the second one is the Einstein telescope with a sensitivity $S_{n}^{1/2}\sim 10^{-24}\,\mathrm{[\rm Hz]}^{-1/2}$ at similar frequency band [75].
Our results show that for stars of about $1.4\,M_{\odot}$ the $f$-mode frequency is near to $2.0$ kHz and the damping time is about $164.0$ ms. Also, as can be seen, at a distance of $10$ kpc, the minimum energy emitted in the GW must be $1.53\times 10^{-7}M_{\odot}$. As can be seen, this required energy is much lower than the energy released in a CCSN explosion, which is about ($10^{-5}-10^{-6}M_{\odot}$ ).
For a fixed central energy density $\rho_{c}=600\,[\rm MeV/fm^{3}]$, varying the anisotropy from $\alpha=-1.0$ to $\alpha=+1.0$ produces measurable shifts in the $f$-mode properties and the GW energy needed for detectability. For $\alpha=-1.0$ we obtain $M=1.1918\,M_{\odot}$, $f=2.1073\,[\rm kHz]$, $\tau=187.033\,[\rm ms]$, with detection thresholds $E^{LV}_{\mathrm{gw}}>1.5409\times 10^{-7}M_{\odot}$ (Advanced LIGO/Virgo, S/N>5>5) and $E^{ET}_{\mathrm{gw}}>3.8523\times 10^{-10}M_{\odot}$ (Einstein Telescope, S/N>5>5). For $\alpha=+1.0$ we find $M=1.7802\,M_{\odot}$, $f=2.0781\,[\rm kHz]$, $\tau=153.927\,[\rm ms]$, and thresholds $E^{LV}_{\mathrm{gw}}>1.4985\times 10^{-7}M_{\odot}$ and $E^{ET}_{\mathrm{gw}}>3.7462\times 10^{-10}M_{\odot}$. This systematic response of $(f,\tau,E_{\mathrm{gw}})$ to $\alpha$-combined with independent constraints on $M$ and $R$-provides a practical diagnostic for future GW asteroseismology: precise measurements can bound the level of anisotropy and, given the higher compactness and typically higher $f$-mode frequencies expected for strange-matter stars, help discriminate hadronic from strange-matter compositions.
@@ 4.4 Tidal deformability
Figure 3: left: Dimensionless tidal deformability against the total mass for different values of $\alpha$. right: Oscillation frequency $f_{f}$ as a function of the dimensionless tidal deformability for five values of $\alpha$. In both panels, the vertical dashed line depicts dimensionless tidal deformability $\Lambda_{1.4}=190^{+390}_{-120}$ from the event GW170817170817 reported by LVC in Ref. [76].
Fig. 3 plots, in its left and right panels respectively, the dimensionless tidal deformability $\Lambda$ against the total mass $M/M_{\odot}$ and the oscillation frequencies of the fundamental mode $f_{f}$ as a function of the dimensionless tidal deformability $\Lambda$ for several choices of the anisotropy parameter $\alpha$. These curves are compared with the observational range for $\Lambda_{1.4}=190^{+390}_{-120}$ reported by LIGO-Virgo [76]. On the left panel, in all cases, the tidal deformability declines steadily as the total mass increases. Additionally, the influence of anisotropy on the deformability response is clearly seen: when $\alpha$ is positive, the calculated $\Lambda$ values are larger for a given mass, whereas negative $\alpha$ leads to smaller values. Importantly, all curves remain within the $\Lambda_{1.4}$ band determined by LIGO-Virgo. On the right panel, the plots show that, as the tidal deformability increases, the $f$-mode frequency decreases monotonically.
@@
The stellar masses considered fall within the intervals $1.36\leq M_{1}/M_{\odot}\leq 1.60$ and $1.17\leq M_{2}/M_{\odot}\leq 1.36$, respectively. In addition, the $50\%$ and $90\%$ credibility contours for the GW170817170817 event, provided by LVC under the low-spin assumption, are included in the plot. For both $\alpha>0$ and $\alpha<0$, the effect of the anisotropy parameter on the tidal deformability becomes evident. Furthermore, we can observe that larger positive values of $\alpha$ help to have dimensionless tidal deformability values closer to the confidence intervals reported in [1].
@@ 5 Conclusions
This article reports the effects of anisotropy on both the stellar structure and $f$ oscillation modes of strange stars. For this aim, we used the stellar equilibrium equations, including the anisotropic factor, and we derived the nonradial oscillation equations involving such a factor within the fully general relativity context. For the fluid contained in the star, we assume that it follows the MIT bag model with vector coupling, and the anisotropy is described by the profile $\sigma=\alpha p_{r}(1-\frac{1}{g_{11}})$.
Regarding the equilibrium configurations, we found that the strange star is affected by anisotropy. For certain ranges of central energy density, we observed both higher (lower) mass and radius values when using higher (lower) values of $\alpha$; this is consistent with what has been reported in the literature for strange stars using the MIT bag model EOS. For $\alpha>0$, increasing $\alpha$ favors the proposed model, bringing the masses closer to the observational data. This suggests that anisotropy could play a relevant role in identifying strange stars.
@@
Finally, we also examine the consistency between the dimensionless tidal deformability of anisotropic strange stars and the observational constraints reported by the LVC for the GW170817170817 event. Our analysis shows that the results presented in this work fall within the observational bounds provided by LVC. Our findings indicate that the tidal deformability increases for positive values of $\alpha$ and decreases for negative values (-$\alpha$). Moreover, several studies in the literature have explored the tidal deformability of strange stars under various theoretical frameworks. For instance, [51] investigates this parameter considering quark matter in the color-flavor-locked (CFL) phase of color superconductivity, [77] examines the role of isospin effects in strange quark matter, and [78] analyzes the tidal deformability assuming a quasiparticle model that incorporates nonperturbative aspects of quantum chromodynamics in the low-density regime. In all these works, as in the present study, constraints from the GW170817170817 event are used to place limits on the properties of strange stars within the corresponding theoretical scenarios.
@@ Acknowledgments
JDVA thanks Universidad Privada del Norte and Universidad Nacional Mayor de San Marcos for the financial support - RR No. 005753\,005753-$2021$-R//UNMSM under the project number B2113178121131781. JMZP acknowledges support from FAPERJ, Process SEI-260003/000308/2024. CVF acknowledges the financial support of the productivity program of the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), with Project No. 304569/2022-4. CHL is thankful to the São Paulo Research Foundation FAPESP (Grant No. 2020/05238-9) and to CNPq (Grants No.401565/2023-8 and 305327/2023-2).
@@ A.2 Perturbation equations
Using the baryon conservation equation, $\nabla_{\mu}(Nu^{\mu})=0$, the local law of energy conservation ($u^{\nu}\nabla^{\mu}T_{\mu\nu}$=0) in terms of Lagrangian perturbations yield:
$\Delta\rho=(\rho+p_{r}+\sigma)\frac{\Delta N}{N}.$ (A.12)
@@
$\hskip-28.45274pt\delta\rho=-(p_{r}+\rho+\sigma)\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]Y_{\ell m}-\rho^{\prime}\frac{e^{-\Lambda}}{r^{2}}WY_{\ell m}.$ (A.13)
@@
$H^{\prime}+\frac{He^{2\Lambda}}{r}\left[4\pi r^{2}(p_{r}-\rho+\sigma)+1+\frac{\ell(\ell+1)}{2}\right]=rK^{\prime\prime}-\frac{8\pi\rho^{\prime}e^{\Lambda}W}{r}-\frac{8\pi e^{2\Lambda}V\ell(\ell+1)}{r}\left(p_{r}\right$.
$\left.+\rho+\sigma\right)-\frac{8\pi e^{\Lambda}W^{\prime}}{r}\left(p_{r}+\rho+\sigma\right)-\frac{Ke^{2\Lambda}}{r}\left[8\pi r^{2}(p_{r}+\rho+\sigma)-1+\frac{\ell(\ell+1)}{2}\right]-e^{2\Lambda}\left[4\pi\rho r^{2}\right$.
@@
$\hskip-11.38092pt-\frac{2K^{\prime}e^{2\Psi}}{r}\left(1-\frac{m}{r}+2\pi r^{2}(p_{r}-\rho)\right)-\frac{e^{2\Psi}H}{r^{2}}\left(4\pi(p_{r}+\rho-\sigma+\Gamma p_{r})r^{2}-2+\frac{4m}{r}+4\pi\sigma r^{2}\frac{dp_{r}}{d\rho}\right)$
$\hskip-11.38092pt+\frac{W^{\prime}e^{2\Psi-\Lambda}}{r^{2}}\left(8\pi\left(\rho+p_{r}+\sigma-\Gamma p_{r}\right)-8\pi\sigma\frac{dp_{r}}{d\rho}\right)+\frac{Ve^{2\Psi}}{r^{2}}8\pi\ell(\ell+1)\left(\rho+p_{r}+\sigma-\Gamma p_{r}-\sigma\frac{dp_{r}}{d\rho}\right)$
$+W\left(\rho^{\prime}-p_{r}^{\prime}\right)\frac{8\pi e^{2\Psi-\Lambda}}{r^{2}}=0,$ (A.17)
@@
$\hskip-11.38092pt-\frac{H^{\prime}}{2}(p_{r}+\rho)e^{\Psi}-e^{\Psi}\Psi^{\prime}(p_{r}+\rho)\left[\frac{H}{2}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-e^{\Psi}\Psi^{\prime}\sigma\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}\right$.
$\hskip-11.38092pt\left.+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-\left\{e^{\Psi}\left(p_{r}+\rho+\sigma\right)\frac{dp_{r}}{d\rho}\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]\right\}^{\prime}-K^{\prime}\sigma e^{\Psi}$
$\hskip-11.38092pt+(p_{r}+\rho)\frac{e^{\Lambda-\Psi}}{r^{2}}{\ddot{W}}+(p_{r}+\rho)e^{\Psi}\left[H^{\prime}-K^{\prime}+\frac{H}{r}e^{2\Lambda}\left(\frac{2m}{r}+8\pi r^{2}p_{r}\right)\right]+(p_{r}+\rho)\left(\Psi^{\prime 2}+\Psi^{\prime\prime}\right$.
$\hskip-11.38092pt\left.-\Psi^{\prime}\Lambda^{\prime}\right)\frac{e^{\Psi-\Lambda}}{r^{2}}W-2(p_{r}+\rho+\sigma)\Psi^{\prime}\frac{e^{\Psi-\Lambda}}{r^{3}}W+\frac{2}{r^{3}}\left(\sigma\Lambda^{\prime}+\frac{3\sigma}{r}-\sigma^{\prime}\right)e^{\Psi-\Lambda}W-\frac{2\sigma}{r^{3}}e^{\Psi-\Lambda}W^{\prime}$
@@
$\hskip-11.38092pt(p_{r}+\rho+\sigma)\frac{\ddot{V}}{e^{2\Psi}}+\frac{H}{2}\left(p_{r}+\rho+\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+\frac{W^{\prime}}{r^{2}e^{\Lambda}}\left(\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+\frac{V\ell(\ell+1)}{r^{2}}\left(\Gamma p_{r}\right$.
$\hskip-11.38092pt\left.+\sigma\frac{dp_{r}}{d\rho}\right)+K\left(\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+p_{r}^{\prime}\frac{W}{r^{2}e^{\Lambda}}-\delta{\tilde{\sigma}}=0,$ (A.20)
where $\delta\sigma=\delta{\tilde{\sigma}}Y_{\ell m}$. It is important to mention that the term $\delta{\tilde{\sigma}}$ is renamed as $\delta{\sigma}$. In addition, for the anisotropic profile $\sigma=\sigma(p_{r},g_{11})$, we found that
@@
This last relation is also used to determine Eqs. (2.17)-(2.22).

### v2

@@ 4.3 On the detectability of the fundamental mode signal
1.0-1.0 $445.450$ $0.8582$ $2.1454$
0.5-0.5 $419.308$ $0.8775$ $2.2194$
$400$ $0.0$ $394.512$ $0.8974$ $2.2435$
+0.5+0.5 $371.023$ $0.9183$ $2.2958$
+1.0+1.0 $348.620$ $0.9399$ $2.3498$
1.0-1.0 $225.026$ $1.2347$ $3.0868$
0.5-0.5 $192.603$ $1.3079$ $3.2698$
$600$ $0.0$ $164.723$ $1.3865$ $3.4663$
+0.5+0.5 $140.722$ $1.4705$ $3.3676$
+1.0+1.0 $120.055$ $1.5576$ $3.8941$
Table 3: The damping time $\tau$ for neutron stars, the energy required to excite the fundamental mode, $E^{LV}_{gw}$ and $E^{ET}_{gw}$, for the signal to be detectable by Advanced LIGO-Virgo and the Einstein Telescope, respectively. For this result, we considered the source at $d\sim 10$ $\rm[\rm kpc]$, a signal-to-noise ratio S/N $\sim 5$, two different central densities, and different values of $\alpha$.
In Table 3, we show the minimum energy required for the fundamental mode to be detected. For this objective, we consider two detectors; the first one is Advanced LIGO-Virgo with a sensitivity of $S_{n}^{1/2}\sim 2\times 10^{-23}\,\mathrm{[\rm Hz]}^{-1/2}$ at $\sim[\rm kHz]$ [10], and the second one is the Einstein telescope with a sensitivity $S_{n}^{1/2}\sim 10^{-24}\,\mathrm{[\rm Hz]}^{-1/2}$ at a similar frequency band [1].
For a fixed central energy density $\rho_{c}=600\,[\rm MeV/fm^{3}]$, varying the anisotropy from $\alpha=-1.0$ to $\alpha=+1.0$ produces measurable shifts in the $f$-mode properties and the GW energy needed for detectability. For $\alpha=-1.0$ we obtain $M=1.5498\,M_{\odot}$, $f=1.8863\,[\rm kHz]$, $\tau=225.026\,[\rm ms]$, with detection thresholds $E^{LV}_{\mathrm{gw}}>1.2347\times 10^{-7}M_{\odot}$ (Advanced LIGO/Virgo, S/N>5>5) and $E^{ET}_{\mathrm{gw}}>3.0868\times 10^{-10}M_{\odot}$ (Einstein Telescope, S/N>5>5). For $\alpha=+1.0$ we find $M=1.6902\,M_{\odot}$, $f=2.1187\,[\rm kHz]$, $\tau=120.055\,[\rm ms]$, and thresholds $E^{LV}_{\mathrm{gw}}>1.5576\times 10^{-7}M_{\odot}$ and $E^{ET}_{\mathrm{gw}}>3.8941\times 10^{-10}M_{\odot}$. This systematic response of $(f,\tau,E_{\mathrm{gw}})$ to $\alpha$-combined with independent constraints on $M$ and $R$-provides a practical diagnostic for future GW asteroseismology: precise measurements can bound the level of anisotropy and, given the higher compactness and typically higher $f$-mode frequencies expected for neutroon stars, help discriminate neutron from strange-matter compositions.
@@ 4.4 Tidal deformability
Figure 3: Left: Dimensionless tidal deformability against the total mass for different values of $\alpha$. Right: Oscillation frequency $f_{f}$ as a function of the dimensionless tidal deformability for nine values of $\alpha$. In both panels, the vertical dashed line depicts dimensionless tidal deformability $\Lambda_{1.4}=190^{+390}_{-120}$ from the event GW170817170817 reported by LVC in Ref. [2].
Fig. 3 plots, in its left and right panels respectively, the dimensionless tidal deformability $\Lambda$ against the total mass $M/M_{\odot}$ and the oscillation frequencies of the fundamental mode $f_{f}$ as a function of the dimensionless tidal deformability $\Lambda$ for several choices of the anisotropy parameter $\alpha$. These curves are compared with the observational range for $\Lambda_{1.4}=190^{+390}_{-120}$ reported by LIGO-Virgo [2]. In the left panel, the tidal deformability decreases monotonically as the total mass increases, up to the maximum-mass configuration. After this point, the curves bend clockwise, and the tidal deformability continues to decrease as the total mass diminishes. Additionally, the influence of anisotropy on the deformability response is clearly seen: when $\alpha$ is positive, the calculated $\Lambda$ values are lower for a given mass, whereas negative $\alpha$ leads to larger values. Importantly, all curves remain within the $\Lambda_{1.4}$ band determined by LIGO-Virgo. On the right panel, the plots show that, as the tidal deformability increases, the $f$-mode frequency decreases monotonically.
@@
The stellar masses considered fall within the intervals $1.36\leq M_{1}/M_{\odot}\leq 1.60$ and $1.17\leq M_{2}/M_{\odot}\leq 1.36$, respectively. In addition, the $50\%$ and $90\%$ credibility contours for the GW170817170817 event, provided by LVC under the low-spin assumption, are included in the plot. For both $\alpha>0$ and $\alpha<0$, the effect of the anisotropy parameter on the tidal deformability becomes evident.
@@ 5 Conclusions
This article reports the effects of anisotropy on both the stellar structure and $f$ oscillation modes of neutron stars. For this aim, we used the stellar equilibrium equations, including the anisotropic factor, and we derived the nonradial oscillation equations involving such a factor within the fully general relativity context. For the fluid contained in the star, we assume that the stellar matter obeys an equation of state derived from a consistent matching between microscopic nuclear theory and perturbative QCD calculations, joined through a piecewise polytropic interpolation scheme, and the anisotropy is described by the profile $\sigma=\alpha p_{r}(1-\frac{1}{g_{11}})^{2}$ (see Ref. [49, 40, 84]).
Regarding the equilibrium configurations, we found that the neutron star is affected by anisotropy. For certain ranges of central energy density, we observed both higher (smaller) mass and radius values when using higher (lower) values of $\alpha$. In addition, the change of $\alpha$ favors the proposed model, bringing the masses closer to the observational data. This suggests that anisotropy could play a relevant role in identifying neutron stars.
@@
Finally, we also examine the consistency between the dimensionless tidal deformability of anisotropic neutron stars and the observational constraints reported by the LVC for the GW170817170817 event. Our analysis shows that the results presented in this work fall within the observational bounds provided by LVC. Our findings indicate that, when $\alpha>0$, the tidal deformability decreases for larger values of $\alpha$ and, when $\alpha<0$, the tidal deformability increases for lower values of $\alpha$. Moreover, several studies in the literature have explored the tidal deformability of compact stars under various theoretical frameworks. For instance, [52] investigates this parameter considering quark matter in the color-flavor-locked phase of color superconductivity, [50] examines the role of isospin effects in strange quark matter, [82] analyzes the tidal deformability assuming a quasiparticle model that incorporates nonperturbative aspects of quantum chromodynamics in the low-density regime, and in [55] to constrain the nuclear matter equation of state from the neutron star tidal deformability. In all these works, as in the present study, constraints from the GW170817170817 event are used to place limits on the properties of compact stars within the corresponding theoretical scenarios.
@@ Acknowledgments
The authors thank Shu Yan Lau, Siddarth Ajith, Victor Guedes, and Kent Yagi for useful discussions regarding the work. JDVA thanks Universidad Privada del Norte and Universidad Nacional Mayor de San Marcos for the financial support - RR No. 005753\,005753-$2021$-R//UNMSM under the project number B2113178121131781. JMZP acknowledges support from FAPERJ, Process SEI-260003/000308/2024. CVF acknowledges the financial support of the productivity program of the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), with Project No. 304569/2022-4. CHL is thankful to the São Paulo Research Foundation FAPESP (Grant No. 2020/05238-9) and to CNPq (Grants No.401565/2023-8 and 305327/2023-2).
@@ A.2 Perturbation equations
Using the baryon conservation equation, $\nabla_{\mu}(Nu^{\mu})=0$, the local law of energy conservation ($u^{\nu}\nabla^{\mu}T_{\mu\nu}$=0) in terms of Lagrangian perturbations yield (see [49]):
$\Delta\rho=(\rho+p_{r})\frac{\Delta N}{N}-\sigma\left(\Delta U^{\theta}_{\theta}+\Delta U^{\phi}_{\phi}\right),$ (A.12)
$\Delta U^{\theta}_{\theta}+\Delta U^{\phi}_{\phi}=\left(K+2e^{-\Lambda}\frac{W}{r^{2}}+\frac{\ell(\ell+1)}{r^{2}}V\right)Y_{\ell m}.$ (A.13)
@@
$\hskip-28.45274pt\delta\rho=-(p_{r}+\rho)\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]Y_{\ell m}-\sigma\left(K+\frac{V}{r^{2}}\ell(\ell+1)\right$.
$\left.+\frac{2e^{-\Lambda}}{r^{3}}W\right)Y_{\ell m}-\rho^{\prime}\frac{e^{-\Lambda}}{r^{2}}WY_{\ell m}.$ (A.14)
@@
$\hskip-11.38092ptH^{\prime}+\frac{He^{2\Lambda}}{r}\left[4\pi r^{2}(p_{r}-\rho)+1+\frac{\ell(\ell+1)}{2}\right]=rK^{\prime\prime}-\frac{8\pi e^{\Lambda}W}{r}\left[\rho^{\prime}+\frac{2\sigma}{r}\right]-\frac{8\pi e^{2\Lambda}V\ell(\ell+1)}{r}\left(p_{r}\right$.
$\hskip-11.38092pt\left.+\rho+\sigma\right)-\frac{8\pi e^{\Lambda}W^{\prime}}{r}\left(p_{r}+\rho\right)-\frac{Ke^{2\Lambda}}{r}\left[8\pi r^{2}(p_{r}+\rho+\sigma)-1+\frac{\ell(\ell+1)}{2}\right]-e^{2\Lambda}\left[4\pi\rho r^{2}\right$.
@@
$\hskip-11.38092pt-\frac{2K^{\prime}e^{2\Psi}}{r}\left(1-\frac{m}{r}+2\pi r^{2}(p_{r}-\rho)\right)-\frac{e^{2\Psi}H}{r^{2}}\left(4\pi(p_{r}+\rho+\Gamma p_{r})r^{2}-2+\frac{4m}{r}\right)$
$\hskip-11.38092pt+\frac{W^{\prime}e^{2\Psi-\Lambda}}{r^{2}}8\pi\left(\rho+p_{r}-\Gamma p_{r}\right)+\frac{Ve^{2\Psi}}{r^{2}}8\pi\ell(\ell+1)\left(\rho+p_{r}+\sigma-\Gamma p_{r}-\sigma\frac{dp_{r}}{d\rho}\right)$
$+W\left(\rho^{\prime}-p_{r}^{\prime}+\frac{2\sigma}{r}\left(1-\frac{dp_{r}}{d\rho}\right)\right)\frac{8\pi e^{2\Psi-\Lambda}}{r^{2}}=0,$ (A.18)
@@
$\hskip-11.38092pt-\frac{H^{\prime}}{2}(p_{r}+\rho)e^{\Psi}-e^{\Psi}\Psi^{\prime}(p_{r}+\rho)\left[\frac{H}{2}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-e^{\Psi}\Psi^{\prime}\sigma\left[\frac{2e^{-\Lambda}}{r^{3}}W\right$.
$\hskip-11.38092pt\left.+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-\left\{e^{\Psi}\left(p_{r}+\rho\right)\frac{dp_{r}}{d\rho}\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]\right$.
$\left.+e^{\Psi}\sigma\frac{dp_{r}}{d\rho}\left(K+\frac{\ell(\ell+1)V}{r^{2}}+\frac{2e^{-\Lambda}W}{r^{3}}\right)\right\}^{\prime}-K^{\prime}\sigma e^{\Psi}+(p_{r}+\rho)\frac{e^{\Lambda-\Psi}}{r^{2}}{\ddot{W}}$
$\hskip-11.38092pt+(p_{r}+\rho)e^{\Psi}\left[H^{\prime}-K^{\prime}+\frac{H}{r}e^{2\Lambda}\left(\frac{2m}{r}+8\pi r^{2}p_{r}\right)\right]+(p_{r}+\rho)\left(\Psi^{\prime 2}+\Psi^{\prime\prime}-\Psi^{\prime}\Lambda^{\prime}\right$.
$\hskip-11.38092pt\left.-\frac{2\Psi^{\prime}}{r}\right)\frac{e^{\Psi-\Lambda}}{r^{2}}W-\frac{2}{r^{3}}\left(\frac{3\sigma}{r}-\sigma^{\prime}-\sigma\left(\frac{2me^{2\Lambda}}{r^{2}}+4\pi e^{2\Lambda}(p_{r}-\rho)r\right)\right)e^{\Psi-\Lambda}W-\frac{2\sigma}{r^{3}}e^{\Psi-\Lambda}W^{\prime}$
@@
$\hskip-11.38092pt(p_{r}+\rho+\sigma)\frac{\ddot{V}}{e^{2\Psi}}+\frac{H}{2}\left(p_{r}+\rho+\Gamma p_{r}\right)+\frac{W^{\prime}}{r^{2}e^{\Lambda}}\Gamma p_{r}+\frac{V\ell(\ell+1)}{r^{2}}\left(\Gamma p_{r}\right$.
$\hskip-11.38092pt\left.+\sigma\frac{dp_{r}}{d\rho}\right)+K\left(\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+W\left(\frac{p_{r}^{\prime}}{r^{2}e^{\Lambda}}+\frac{2\sigma}{e^{\Lambda}r^{3}}\frac{dp_{r}}{d\rho}\right)-\delta{\tilde{\sigma}}=0,$ (A.21)
where $\Gamma=(p_{r}+\rho)\frac{dp_{r}}{d\rho}$ and $\delta\sigma=\delta{\tilde{\sigma}}Y_{\ell m}$. It is important to mention that the term $\delta{\tilde{\sigma}}$ is renamed as $\delta{\sigma}$. In addition, for the anisotropic profile $\sigma=\sigma(p_{r},g_{11})$, we found that
@@
This last relation is also used to determine Eqs. (2.17)-(2.22). It is important to mention that, in the Cowling approximation, $h_{\mu\nu}$=0, Eqs. (A.14)-(A.22) reproduce the nonradial oscillation equations reported in Ref. [29].